The nut store sells walnuts for 4.00 a pound and cashews for 7.00 a pound how many pounds of cashews should be mixed with 10 pounds of walnuts to obtain a mixture that sells for 5.50 a pound?
Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE WHICH WILL HELP YOU TO GET AN INSIGHT INTO THESE PROBLEMS AND THE METHOD OF APPROACH TO SOLVE THEM
--------------------------------------------------------------------------
How much pure alcohol must a nurse add to 8cc of a 70% solution to strengthen it to a 90% solution? Just need to know if this is right and if not what is right.
.90(8)+ 0(x) = .70(x+8)....NO ...FIRST WRITE DOWN WHAT IS X?I THINK X IS QTY.OF
PURE ALCOHOL TO BE ADDED TO 8 CC OF 70% ALCOHOL TO GET X+8 CC OF 90% SOLUTION.SO THE EQN. SHOULD BE X*100/100+8*70/100=(X+8)*90/100
7.2 = .70x+5.6.....X+5.6=0.9(X+8)
1.6 = 70x.......X-0.9X=7.2-5.6=1.6
2.29c = x.......0.1X=1.6
..................OR X=1.6/0.1=16 CC.
SEE THE FOLLOWING EXAMPLE WHICH WILL HELP YOU TO GET AN INSIGHT INTO THESE PROBLEMS AND THE METHOD OF APPROACH TO SOLVE THEM
------------------------------------------------------------------------------
answer 15572 by venugopalramana(1008)on 2006-02-26 07:15:31 
PLEASE HELP ASAP: Ziggy's famous yogurt blends regular yogurt that is 3% fat with its no fat yogurt to obtain low fat yogurt that is 1% fat. How many pounds of regular and how many pounds of non-fat yogurt should be mixed to obtain 60 pounds of lowfat yogurt.
PLEASE HELP ASAP. thank you
THESE ARE MATERIAL BALANCE PROBLEMS.THE PRINCIPLE IS TO APPLY
TOTAL OF ALL INPUTS =TOTAL OF ALL OUTPUTS..
THIS PRINCIPLE CAN BE APPLIED TO TOTAL MIXTURE AS A WHOLE AS WELL AS INDIVIDUAL COMPONENTS OF THE MIXTURE.LET US SEE THE APPLICATION USING YOUR PROBLEM.
HERE THE MIXTURE COMPRISES 2 INPUTS-REGULAR YOGURT (RY) & NO FAT YOGURT (NFY)
AND ONE OUT PUT-LOW FAT YOGURT (LFY).THE COMPONENT OF IMPORTANCE IN THE MIXTURE IS FAT CONTENT.SO WE TAKE 2 BALANCES HERE ..ONE FOR THE TOTAL MIXTURE AND ANOTHER FOR COMPONENT OF FAT IN THE MIXTURE.
I..TOTAL BALANCE...
INPUTS
1.QTY.OF.RY=X POUNDS
2.QTY OF NFY=Y POUNDS
OUT PUT
1.QTY.OF LFY=60 POUNDS
SO APPLYING
TOTAL OF ALL INPUTS =TOTAL OF ALL OUTPUTS.....WE GET
X+Y=60.............................I
II..COMPONENT BALANCE..HERE IT IS FAT .
INPUTS
1.QTY.OF FAT IN RY=X*3/100=3X/100 POUNDS
2.QTY OF FAT IN NFY=Y*0/100=0 POUNDS
OUT PUT
1.QTY.OF FAT IN LFY=60*1/100=60/100 POUNDS
SO APPLYING
TOTAL OF ALL INPUTS =TOTAL OF ALL OUTPUTS.....WE GET
3X/100 + 0=60/100.............................II
3X=60
X=20 POUNDS. OF REGULAR YOGURT
Y=60-20=40 POUNDS OF NO FAT YOGURT.